OpenStudy (anonymous):

The blood value of patient were pH=7.03 and [CO2]=1.1mM. What is the patient's blood[HCO3^-] and how much of the normal [HCO3^-] has been used in buffering the acid causing the condition?

OpenStudy (aaronq):

So you have this equilibrium: \(CO_2+H_2O \rightleftharpoons H_2CO_3+H_2O \rightleftharpoons HCO_3^-+H_3O^+ \) We can ignore the water because it's concentration is much higher than the rest of the substances present. So we have: \(K^1_{eq}=\dfrac{[H_2CO_3]}{[CO_2]}\) and \(K^2_{eq}=\dfrac{[H_3O^+][HCO_3^-]}{[H_2CO_3]}\) We want just 1 equation with what we are given, so, do some algebra. rearrange the first: \([H_2CO_3]=K^1_{eq}*[CO_2]\) sub into second: \(K^2_{eq}=\dfrac{[H_3O^+][HCO_3^-]}{K^1_{eq}*[CO_2]}\) we have: \(K^1_{eq}*K^2_{eq}=\dfrac{[H_3O^+][HCO_3^-]}{[CO_2]}\) You're given pH, so you know \([H_3O^+]\)...so it seems that you can simply say: \([H_3O^+]=[HCO_3^-]\)...But i'm not sure, did anything else happen? What acid are they talking about?